OpenStudy (anonymous):

g(x)= (x^2+x+1)(4-x)/ 2x-1 differentiate the given function

OpenStudy (anonymous):

f(u,v) = u/v f' = (u'v-v'u)/(v^2) does this help ?

OpenStudy (anonymous):

no not really....i understand what you put up, but there's like two part to the numerator

OpenStudy (anonymous):

so you can either use the product rule or just simplify the numerator then differentiate

OpenStudy (anonymous):

Differentiation using the quotient rule

OpenStudy (anonymous):

\[y = \frac{ u }{ v } \rightarrow \frac{ dy }{ dx } = \frac{ v \frac{ du }{ dx }-u \frac{ dv }{dx } }{ v^2 }\]